A ball player wishes to determine pitching speed by throwing a ball horizontally from an elevation of 5 m above the ground. The player sees the ball land 20m down range. What is the player's pitching speed?
1
Expert's answer
2018-09-26T10:18:09-0400
We can find the time of flight from the formula:
h=(gt^2)/2.
t=√(2h/g).
The player's pitching speed is
v=x/t=x/√(2h/g)=x√(g/2h)
v=20√(10/(2(20)))=10 m/s.